Please explain the steps as well; if possible
3. Two masses, approach each other and collide head-on. direction) has a mass which is three times that of the mass traveling to the right. If they both were initially traveling a 5 m/s and stick together after the collision, determine the velocity of the combination after the collision. The mass traveling to the left (i
Answer #1

As per the data

Two Masses approach each other

Mass travelling to rigid directlion (M) =m1

​Mass travelling to right direction (M) =3m1

Velocity of bodies

u1 =5 m/s u2 =5m/s

Applying law of conservation of momentum

m1 u1 + m2 u2 =m1 v1 + m2 v2

As the two masses stick together v1 = v2 =v

Substiting Values

3m1 (5) + (5) m1 = 3 m1 (v) + m 1(v)

15 m1+ 5m1 =v (3m1 + m1)

20 m1 = v ( 4 m1)

Velocity =5 m/s

VELOCITY after the collission = 5 m/s